# Spurious reconstruction from brain activity

### v3

@@ Acknowledgments
We thank our laboratory team, especially Eizaburo Doi, Haibao Wang, Kenya Otsuka, Hideki Izumi, and Matthias Mildenberger for their invaluable comments and suggestions on the manuscript. This work was supported by the Japan Society for the Promotion of Science (JSPS: KAKENHI grants JP20H05954, JP20H05705 to Y.K, JP21K17821 to Y.N, and 22KJ1801 to K.S.), Japan Science and Technology Agency (JST: CREST grants JPMJCR18A5, and JPMJCR22P3 to Y.K.), and New Energy and Industrial Technology Development Organization (NEDO: commissioned project, JPNP20006 to Y.K.).
Supporting information
@@
S1 Fig. UMAP visualization of CLIP text features in the Natural Scene Dataset (NSD) using default parameters. Similar to Fig 4A in the main text, this scatter plot visualizes the distribution of semantic concepts within the NSD. The gray points represent training samples, and the orange points represent test samples. Unlike Fig 4A, which used parameters optimized for clustering visualization, this figure employs default UMAP settings. Despite the difference in parameters, the visualization still reveals distinct clusters and substantial overlap between training and test samples.
@@
S2 Fig. UMAP visualization of CLIP text features in the Deeprecon dataset. This scatter plot visualizes the distribution of semantic concepts within the Deeprecon data, with the gray points representing training samples and the orange points indicating test samples. Unlike the NSD dataset (Fig 4A in the main text), this visualization demonstrates a clearer separation between training and test samples. This distinction highlights the Deeprecon dataset’s intentional design to differentiate object categories between training and test sets.
@@
S3 Fig. Confusion matrices from the hold-out analyses. (A) Clustering results of CLIP text features. This panel shows the outcome of applying $k$-means clustering to the UMAP embedding of CLIP text features from the NSD dataset (Fig 4A in the main text). These clustering results were used to determine the assignment of training and test samples under the hold-out split condition (Fig 6A in the main text). (B) Confusion probability matrices for cluster identification. These matrices illustrate the performance of cluster identification under different conditions. The top row shows results under the hold-out split conditions, while the bottom row presents results under the naive split conditions. For each row, the left matrix represents CLIP text features, and the right matrix represents CLIP vision features. In each matrix, the cell at position $(i,j)$ indicates the proportion of samples originally from cluster $i$ that were predicted to belong to cluster $j$. As visualized in the confusion matrices, the hold-out split condition, where entire semantic clusters are excluded from training, reveals the challenges in zero-shot prediction of novel concepts. In contrast, the naive split condition, where samples from all clusters are included in the training, demonstrates the model’s performance when semantic concepts overlap between training and test sets.
@@
S4 Fig. Extended simulation results for clustered data analysis. (A) Relationship between cluster identification accuracy and cluster variance ratio. Each subplot represents a different number of clusters used in model training. The $x$-axis represents the cluster variance ratio $\sigma_{\mathrm{intra}}^{2}/\sigma_{\mathrm{inter}}^{2}$, and the $y$-axis represents the cluster identification performance. The green lines represent results for in-distribution test samples, and the orange lines show outcomes for out-of-distribution (OOD) samples. As the cluster variance ratio increases, clusters tend to overlap more, leading to difficulty with identification in the in-distribution condition. However, larger cluster variance also allows for more extensive coverage of the feature space, potentially improving OOD prediction. Consequently, OOD curves typically show a peak at intermediate variance ratios, reflecting a balance between cluster distinctiveness and feature space coverage. (B) Heatmap of cluster identification accuracy. These matrices visualize how accuracy changes with varying numbers of training clusters and feature space dimensions. The cluster variance ratio ($\sigma_{\mathrm{intra}}^{2}/\sigma_{\mathrm{inter}}^{2}$) is fixed at $0.1$ for all simulations. The left matrix shows results for in-distribution samples, while the right matrix displays outcomes for out-of-distribution (OOD) samples. In each matrix, cell $(i,j)$ indicates the cluster identification accuracy when the feature dimension is $i$ and the number of clusters used in training is $j$. This view illustrates the interplay between feature dimensionality and the number of training clusters in determining the model’s generalization capabilities. Notably, in the OOD condition, as the feature dimensionality increases, a greater number of training clusters becomes necessary to achieve a robust zero-shot prediction.
@@
S5 Fig. Frame similarity analysis in the dataset of Nishimoto et al. [89]. Movie scenes were detected from the movie stimuli. Then, for the first and last frames of each test scene, we identified the frames from the training movies with the closest Euclidean distance. Our analysis revealed that $37$ out of the $48$ scenes in the test set contained frames that were nearly identical to those in the training set.

### v5

@@ Acknowledgments
We thank our laboratory team, especially Eizaburo Doi, Haibao Wang, Kenya Otsuka, Hideki Izumi, and Matthias Mildenberger, for their invaluable feedback and insightful suggestions on the manuscript.
Appendix A Supplementary figures
Fig. A1: Reconstructions from the NSD dataset using a sample size matched to the deeprecon dataset. The figure follows the format of Fig. 2 in the main text. The text-guided reconstruction and MindEye2 methods did not show a large performance drop, even when the training sample size of the NSD dataset was reduced to match that of the Deeprecon dataset. This result supports that the degraded performance in the Deeprecon is not simply due to its smaller sample size compared to the NSD.
@@ Appendix A Supplementary figures
Fig. A2: UMAP visualization of CLIP text features in the Natural Scene Dataset (NSD) using default parameters. This scatter plot, analogous to Fig. 4A in the main text, depicts the distribution of text features within the NSD. The gray points represent training samples, and the orange points represent test samples. Unlike Fig. 4A, which used the parameters optimized for clustering visualization, this figure employs the default UMAP settings. Despite the absence of parameter tuning, the plot still reveals discernible clusters and considerable overlap between training and test samples, indicating that the observed clustering pattern does not hinge on specialized UMAP configurations.
@@
Fig. A3: UMAP visualization of CLIP text features in the Deeprecon dataset. This scatter plot visualizes the distribution of semantic features within the Deeprecon data, with the gray points representing training samples and the orange points indicating test samples. Unlike the NSD dataset (Fig. 4A in the main text), this visualization demonstrates a clearer separation between training and test samples. This distinction highlights the Deeprecon dataset’s intentional design to differentiate object categories between training and test sets.
Fig. A4: UMAP visualization of the latent features of MindEye2 (OpenCLIP ViT/bigG-14) in the NSD and Deeprecon datasets. The plots follow the format of Fig. 4A and Fig. A3. While these features are extracted from images, not from text captions, the plot still reveals a cluster structure, with considerable overlap between the training and test sets in the NSD dataset, whereas the Deeprecon dataset exhibits noticeably less overlap.
Fig. A5: Distribution of similarity scores between training and test stimuli. Violin plots show the distribution of DreamSim-based similarity scores (Fu et al., 2024) between each dataset’s training set and its test stimuli (green) compared to the similarity distributions between the training set and an independent dataset (orange; randomly selected 1000 images in CC3M; Sharma et al., 2018). For each dataset, we plotted the distribution of similarity scores from the top $5\%$ most similar training images to either test or independent set images. Results remained consistent when varying the percentage or number of selected training images. A large deviation between the test- and independent-set distributions indicates that the test set is biased toward the training set, while smaller deviations suggest greater test set independence. This analysis was performed on seven datasets: Deeprecon ($1,200$ training and $90$ test images) and NSD (8,859 training and $982$ test images), and was further extended to THINGS-fMRI1 ($8,640$ training and $100$ test images; Hebart et al., 2023), THINGS-MEG ($22,248$ training and $200$ test images; Hebart et al., 2023), DreamDiffusion ($1,330$ training and $333$ test images; Bai et al., 2024), Brain2GAN (4,000 training and 200 test images; Dado et al., 2024), and TVSD ($22,248$ training and $100$ test images; Papale et al., 2025). Most datasets exhibited higher training-test similarity compared to training-independent similarity, suggesting these splits are not optimal for evaluating zero-shot prediction. In contrast, the Deeprecon dataset was specifically designed to exclude overlapping training categories, facilitating a more suitable evaluation of generalizability performance. The distributions of training-independent similarity were consistent across the datasets, suggesting comparable levels of representativeness among the datasets.
@@
Fig. A6: Confusion matrices from the hold-out and naive splits. (A) Clustering results of CLIP text features. The $k$-means clustering was applied to the UMAP embedding of CLIP text features from the NSD (Fig. 4A in the main text). These clustering results were used to make a training–test split under the hold-out split condition (Fig. 6A). (B) Confusion probability matrices for cluster identification. Left and right matrices represent CLIP text and vision features, respectively, with each cell $(i,j)$ indicating the proportion of samples from cluster $i$ predicted as cluster $j$. The top row shows results for the hold-out split, where entire semantic clusters are excluded from training, highlighting challenges in zero-shot prediction of novel categories. The bottom row shows the naive split, where semantic clusters appear in both training and test sets, demonstrating the translators’ ability to identify learned semantic categories.
Fig. A7: Reconstruction analysis under hold-out split conditions using the NSD dataset. Stimulus images were clustered via UMAP of CLIP text features into “held-out” and “training” groups; models were trained only on training clusters and evaluated on held-out clusters using independent brain data. (A) DreamSim similarity distributions for various numbers of held-out clusters. The figure shows similarities between the training set and the hold-out test set (green) versus those between the training set and an independent dataset (CC3M; orange) as in Fig. A5. The x-axis indicates the number of excluded clusters. Despite removals, the test set remains more similar to the training set than to CC3M. The “similarity-matched hold-out” condition aligns these similarities with Deeprecon (see Fig. A5). The “naive split” (right side of the dashed line) randomly removes samples from the original training set so that the total number of training samples matches that of the similarity-matched hold-out split. Unlike the similarity-matched hold-out split, the naive split allows cluster overlap between training and test sets. (B, C) Reconstructions under different split conditions for StableDiffusionReconstruction (B) and Brain-Diffuser (C). Each row shows the original test image (top), followed by reconstructions from the naive split, the 1-cluster hold-out, and the similarity-matched hold-out. All splits use the same number of training samples for fair comparison. For StableDiffusionReconstruction, regression parameters were adjusted to accommodate the smaller training set, yet the outputs remained low-quality. Although the naive-split results retain some semantic similarity to the test images, using a 1-cluster hold-out led to visibly and semantically different reconstructions from the original test images. Brain-Diffuser reconstructions reveal clearer differences across conditions: naive-split outputs closely resemble the test images, 1-cluster hold-out reconstructions preserve the overall layout with minor semantic changes, and similarity-matched hold-out reconstructions deviate substantially in both visual and semantic content.
Fig. A8: Recovery check for MindEye2. The panels follow the format of Fig. 7 in the main text. MindEye2 uses a two-stage generation process. Left (SDXL unCLIP): In the initial stage, latent features are fed into a fine-tuned SDXL (Podell et al., 2023) with unCLIP technique (Ramesh et al., 2022; Scotti et al., 2024) to generate images, which closely match the originals. Right (MindEye2): These initial images are then refined in a second stage using base SDXL and captions predicted from the latent features (via GiT Image2Text modules; Wang et al., 2022). Because this second stage incorporates text information, the final MindEye2 outputs show lower fidelity than the SDXL unCLIP results.
Fig. A9: Impact of the nonlinear translator on MindEye2 reconstructions. The first row shows the test images (Deeprecon), and the second row presents the full MindEye2 reconstructions, which use a nonlinear translator followed by a two-stage generator (SDXL unCLIP and then SDXL with predicted captions). To isolate the effects of the generator utilizing GiT Image2Text modules, the third and fourth rows compare only the first-stage outputs (SDXL unCLIP) using a nonlinear translator (third row) versus a linear translator (fourth row). Nonlinear translators often revert to object categories seen in the training set, whereas linear translators do not exhibit this tendency, suggesting that nonlinear models may be more prone to output dimension collapse.
@@
Fig. A10: Extended simulation results for clustered data analysis. (A) Relationship between cluster identification accuracy and cluster variance ratio. Each subplot represents a different number of clusters used in model training. The $x$-axis represents the cluster variance ratio $\sigma_{\mathrm{intra}}^{2}/\sigma_{\mathrm{inter}}^{2}$, and the $y$-axis represents the cluster identification performance. The green and orange lines indicate in-distribution and out-of-distribution (OOD) test samples, respectively. Higher variance ratios increase cluster overlap, making in-distribution identification harder. However, they also expand feature space, improving OOD prediction. OOD performance peaks at intermediate variance ratios, where cluster separability and feature space coverage are balanced. (B) Heatmap of cluster identification accuracy. These matrices visualize how accuracy changes with varying numbers of training clusters and feature space dimensions. The cluster variance ratio ($\sigma_{\mathrm{intra}}^{2}/\sigma_{\mathrm{inter}}^{2}$) is fixed at $0.1$ for all simulations. The left matrix represents in-distribution samples, and the right represents OOD samples. Each cell $(i,j)$ indicates the accuracy for the feature dimension $i$ and $j$ training clusters. In the OOD condition, higher feature dimensionality requires more training clusters for robust zero-shot prediction and this relationship is linear, not exponential.
@@
Fig. A11: Frame similarity analysis in the dataset of (Nishimoto et al., 2011). Movie scenes were detected from the movie stimuli. Then, for the first and last frames of each test scene, we identified the frames from the training movies with the closest Euclidean distance. Our analysis revealed that 37 out of the 48 scenes in the test set contained frames that were nearly identical to those in the training set.